ROTATIONAL SPECTROSCOPIC STUDY OF He_{n}-OCS CLUSTERS

\underline{YUNJIE XU} and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB, Canada, T6G 2G2.

The rotational spectrum of the He_{2}-OCS complex^{a} was reported by us two years ago at this meeting. More recently, Tang and McKellar investigated the infrared spectra of the He_{n}-OCS with n=1 to 8.^{b} In this talk, we present detailed isotopic studies of the series of He_{n}-OCS (n=2-8) clusters that were done using a cavity Fourier transform microwave spectrometer. Nuclear hyperfine structures due to the quadrupolar ^{33}S and ^{17}O nuclei were observed and analyzed. The resulting isotopic rotational, distortion, and hyperfine constants were used to derive information about the structures and dynamics of these clusters. Investigations of microwave spectra of higher He_{n}-OCS clusters are currently underway.

aY. Xu and W. Jäger, Chem. Phys. Lett. 350, 417 (2001).

bJ. Tang and A. R. W. McKellar, private communication (2001).